# references.bib

### v1

@@ 2.4. OS Support for Hardware QoS Enforcement
Management of QoS enforcement schemes is preferably managed through an operating system-level interface. In Linux, this interface is known as the User Interface for Resource Control (resctrl(yu_24_nodate). resctrl enables the definition of resource allocation groups that dictate resource behavior. Logical processor cores or even individual processes may be assigned to a resource group and will subsequently follow the allocation configuration of that group. Resource allocation groups are mapped directly to hardware classes of service (CLOS or COS). Therefore, the number of available resource groups is limited and dependent on the CPU model. When individual processes are assigned to resource groups, the rules of the core executing that process are modified at the time of context switching such that the rules only apply to the correct process.
@@ 2.5. A Note on Shared Resource Monitoring Features
In addition to the enforcement features discussed, the Intel RDT and AMD QoSE feature sets include monitoring support for shared cache and memory bandwidth. These monitoring features are, similarly to their enforcement counterparts, accessible through the resctrl OS interface (yu_24_nodate).
Due to the indirect nature of the mechanisms provided by QoS enforcement features, tuning behavior is highly dependent on the dynamics of the workloads of a system. Thus, actual behavior may be difficult to predict (sohal_closer_2022; farina_assessing_2022; xiang_emba_2019) and it is therefore recommended that actual resource usage is monitored and taken into account by controllers (intel_corporation_intel_2023). Monitoring features therefore play a crucial role in effective feature utilization as they allow controllers to observe changes in metrics resulting from a tuning decision.
@@ 3. Review Motivation
Through observation of published literature, it is clear that QoS enforcement hardware features are conspicuous by their absence in general-purpose cloud computing resource management literature reviews, see Table 1.
@@
(mittal_survey_2017)
(gracioli_survey_2015; lugo_survey_2022)
(kawashima_vision_2021)
(hameed_survey_2016; liaqat_federated_2017; hong_resource_2019; bermejo_virtual_2020; helali_survey_2021; chaurasia_comprehensive_2021; carrion_kubernetes_2022; huang_survey_2023)
@@
Cloud environments rarely see hard execution time deadlines on tasks, as is common in real-time systems, but rather statements of performance intent in the form of service level objectives (SLOs). Resource management optimization for cloud environments is therefore typically discussed in a broader context and at a higher level of abstraction than real-time systems. Thus, lower-level sources of congestion such as LLC and memory bandwidth are often overlooked, excluded, or disregarded in cloud resource management surveys (hameed_survey_2016; liaqat_federated_2017; hong_resource_2019; bermejo_virtual_2020; helali_survey_2021; chaurasia_comprehensive_2021; carrion_kubernetes_2022; huang_survey_2023). Even so, increased resource efficiency and task consolidation achieved through QoS enforcement feature utilization is not exclusive to real-time systems. We thus recognize the need for a review of literature approaching shared resource management by utilizing QoS enforcement features with a general-purpose cloud-computing point of view.
This review serves to gain insight into how the academic community is approaching resource management problems specifically by using hardware-implemented QoS enforcement features, and to identify current problems and opportunities for improvement in general-purpose cloud computing environments.
@@ 3.1. Related Surveys
A few systematic reviews have previously been conducted that cover neighboring and overlapping areas of research. In their (gracioli_survey_2015) article, gracioli_survey_2015 (gracioli_survey_2015) surveyed cache management approaches for embedded real-time systems. They recognize that server processors supporting Intel CAT, the only commercially available feature at the time, are not targeted at the embedded systems market surveyed in their paper. However, they still make the case for cache allocation’s usefulness for the optimization of soft real-time applications.
@@
Publications with a focus on resource management in the cloud often approach the problem at a higher abstraction layer than papers with lower-level performance as their context, such as those from the real-time computing community. This often means that CPU cache and memory bandwidth management on the node level is out of scope. Such is the case in the surveys by hameed_survey_2016 (hameed_survey_2016), liaqat_federated_2017 (liaqat_federated_2017), and hong_resource_2019 (hong_resource_2019).
@@
With a clear focus on cache partitioning, mittal_survey_2017 (mittal_survey_2017) surveyed techniques for mitigating the increased shared cache congestion that follows from increasing core counts in processors. This includes early publications utilizing CAT, as well as a large number of software-based approaches using techniques such as page coloring.
@@
bermejo_virtual_2020 (bermejo_virtual_2020), helali_survey_2021 (helali_survey_2021), and chaurasia_comprehensive_2021 (chaurasia_comprehensive_2021) surveyed techniques of workload consolidation in the cloud. The first with a specific focus on virtual machine (VM) consolidation. kawashima_vision_2021 (kawashima_vision_2021) acknowledge the usefulness of cache partitioning when optimizing the performance of virtual network functions (VNFs) that power much of the networking in modern clouds.
@@
lugo_survey_2022 (lugo_survey_2022) surveyed works on process interference for real-time systems published between 2015 and 2020. The works included proposed techniques that reduce congestion in LLC, main memory, as well as in the main memory bus. A few of the works utilize QoS enforcement features, while others use software-based approaches.
@@
With the rise of Kubernetes as the de-facto standard container orchestrator, carrion_kubernetes_2022 (carrion_kubernetes_2022) identifies the need to survey the current approaches and ongoing issues in Kubernetes scheduling. They find that Kubernetes is used in many different environments, running different types of workloads on different types of hardware, which dictates the type of scheduling approaches leveraged.
@@
In a recent publication, huang_survey_2023 (huang_survey_2023) review works on resource management in the context of cloud-native mobile computing. This has received an increasing amount of attention recently due to the architecture of fifth and sixth-generation mobile communication networks (5G and 6G) being based on cloud-native technology. Most of the works reviewed focus on the management of networking resources through the utilization of network slicing or network virtualization.
@@ 4.1. Search Strategy
Our publication search strategy is based on the process outlined by wohlin_guidelines_2014 (wohlin_guidelines_2014) and follows a simple algorithm. First, a starting set of core publications is identified by querying a digital literature database using a defined set of search terms and inclusion criteria. Then, both backward and forward snowballing is used to identify additional papers that fit the inclusion criteria. The snowballing process is repeated after identified papers are added to the set of included papers. This process is repeated until no additional papers that fit the inclusion criteria are found. This process is outlined in Figure 5.
Figure 5. Reference snowballing search strategy used.
@@
Our motivations for selecting this approach instead of a more traditional database search approach (i.e. kitchenham_evidence-based_2004 (kitchenham_evidence-based_2004)) are as follows: We survey how the academic community leverages a defined set of hardware features in research. We can thus find our starting set of articles by explicitly searching for the names of the features. This should prove a good basis for snowballing as authors typically cite a small set of early papers as sources for further details of a feature. Such publications become natural “roots” in our reference tree. Moreover, it was shown by badampudi_experiences_2015 (badampudi_experiences_2015) that snowballing is comparable, if not more reliable, than a traditional database search approach.
@@ 4.2. Threats to Review Validity
The snowballing process is highly dependent on the quality of the starting set of papers used to seed the search. A starting set that is too small, or not diverse enough can leave “dark spots” where many papers are left out. One reason for this is a same author bias that may emerge. As researchers often cite their previous work due to relevance, a snowballing search may become heavily biased towards a few researchers’ works (jalali_systematic_2012). The same phenomenon may be true for publishers instead of individual researchers (wohlin_guidelines_2014).
@@ 4.3. Inclusion Criteria
No starting date of the publication time span is defined as hardware QoS features have only recently become available in commodity off-the-shelf server processors. Intel’s CAT was originally introduced in the Xeon E5-2600 v3 family of processors (herdrich_cache_2016) which launched in the third quarter of 2014. Thus, 2014 serves as a starting date for CAT publications.
We limit our review to hardware-based control mechanisms even though software-based QoS enforcement approaches have existed for many years (mittal_survey_2017). Software-based approaches such as page coloring (taylor_tlb_1990) or MemGuard (yun_memguard_2013) provide similar functionality to their hardware-based counterparts and have likely inspired the design of the features now implemented in hardware. However, the native and purpose-built nature of hardware-level control mechanisms means their utilization is preferable in most cases, which has started to make software-based approaches obsolete. Additionally, the implementation specifics of hardware-level control mechanisms mean their areas of applicability and their limitations may vary from software-level approaches.
@@ 4.4. Starting Set Identification
Google Scholar [^3] http://scholar.google.com was used as a search engine when querying for potential starting set articles. This had the advantage of minimizing potential publisher bias as the search covers a large number of sources. A separate search was conducted for each of the hardware features of interest. The exact search terms used were:
@@
A set of seed publications were selected from these lists of publications based on our inclusion criteria, publisher diversity, and relevance. This start set of papers are listed in Table 2 and their relationship graph may be seen in Figure 6.
@@
(funaro_ginseng_2016) 1 funaro_ginseng_2016
(xu_vcat_2017) 1 xu_vcat_2017
(selfa_application_2017) 1 selfa_application_2017
(xu_dcat_2018) 1 xu_dcat_2018
(pons_improving_2018) 1 pons_improving_2018
(park_hypart_2018) 2 park_hypart_2018
(xiang_dcaps_2018) 1 xiang_dcaps_2018
(kim_application_2019) 1 kim_application_2019
(park_copart_2019) 2 park_copart_2019
(xiang_emba_2019) 2 xiang_emba_2019
(zhang_libra_2021) 3 zhang_libra_2021
(sohal_closer_2022) 1 sohal_closer_2022
(farina_assessing_2022) 2 farina_assessing_2022
@@
Figure 6. Arc diagram illustrating the reference relationships between the papers selected for the starter set. Publications are chronologically ordered from top to bottom. The size of the nodes corresponds to the number of connections they have.
@@ 4.5. Reference Snowballing
Using the snowballing search method, we iterated upon the defined starting set. At the time of the final snowballing algorithm iteration, the search set had increased to $3685$ references. Out of these $3685$ publications, $79$ papers were identified to meet the inclusion criteria of our review. Figure 7 illustrates the included articles’ publication dates and how interest from researchers in both academia and industry has evolved over time.
@@
Figure 7. Distribution of articles utilizing platform features grouped by their year of publication. Papers are classified as academic when all authors are affiliated with academic institutions, whereas cooperative publications have one or more authors affiliated with a private company. Industry papers have no affiliation with an academic institution. (*2023 is limited to the first six months of the year.)
@@ 5. Taxonomy
Figure 8. Taxonomy organization.
@@ 5.1. Feature Evaluations
The function of QoS enforcement features implemented in hardware has been experimentally analyzed by a number of different works (sohal_closer_2022; zhang_libra_2021; herdrich_cache_2016; farina_assessing_2022; ihre_sherif_evaluation_2021; schramm_impact_2019; kim_simple_2017; veitch_performance_2017; castiglione_optimal_2018). It is important to verify the behavior of hardware features so that they may be used as a basis for other research. However, the functionality may change between chip generations due to architectural differences (sohal_closer_2022). This further highlights the importance of analyzing the behavior of new implementations as the features mature over time. Additionally, some works (bechtel_denial--service_2022; qiu_is_2021; farshin_reexamining_2020) evaluate if QoS enforcement features may be used as solutions to problems in other research areas. Other publications (shang_coplace_2021) use QoS enforcement features to demonstrate problems and subsequent solutions to resource management problems.
@@
(herdrich_cache_2016) Introduction and feature showcase of CAT.
(sohal_closer_2022) Experimental evaluation of CAT and MBA.
(zhang_libra_2021) Highlights the shortcomings of MBA.
(farina_assessing_2022) Experimental evaluation of MBA for RT systems.
(ihre_sherif_evaluation_2021) Experimental evaluation of CAT.
(schramm_impact_2019) Experimental evaluation of CAT for VNF systems.
(qin_when_2018) Study of cache partitioning in correlation to workload cache characteristics.
(kim_simple_2017) Evaluation of cache partitioning as a tool to meet workload latency targets.
(veitch_performance_2017) Experimental evaluation of CAT for VNF systems.
(castiglione_optimal_2018) Highlights the potential usefulness of CAT for security purposes.
@@
(qiu_is_2021) Evaluation of FaaS for LC workload hosting.
(bechtel_denial--service_2022) Evaluation of LLC partitioning as DoS protection on iGPU platforms.
(farshin_reexamining_2020) Evaluation of Intel Data Direct I/O.
(shang_coplace_2021) Intelligent Guest to Host Page Mapping.

### v2

@@ 2.5. OS Support for Hardware QoS Enforcement
Management of QoS enforcement schemes is preferably managed through an operating system-level interface. In Linux, this interface is known as the User Interface for Resource Control (resctrl(yu_24_nodate, 116). resctrl enables the definition of resource allocation groups that dictate resource behavior. Logical processor cores or even individual processes may be assigned to a resource group and will subsequently follow the allocation configuration of that group. Resource allocation groups are mapped directly to hardware classes of service (CLOS or COS). Therefore, the number of available resource groups is limited and dependent on the CPU model. When individual processes are assigned to resource groups, the rules of the core executing that process are modified at the time of context switching such that the rules only apply to the correct process.
@@ 2.6. A Note on Shared Resource Monitoring Features
In addition to the enforcement features discussed, the Intel RDT and AMD QoSE feature sets include monitoring support for shared cache and memory bandwidth. These monitoring features are, similarly to their enforcement counterparts, accessible through the resctrl OS interface (yu_24_nodate, 116).
Due to the indirect nature of the mechanisms provided by QoS enforcement features, tuning behavior is highly dependent on the dynamics of the workloads of a system. Thus, actual behavior may be difficult to predict (sohal_closer_2022, 95, 27, 109) and it is therefore recommended that actual resource usage is monitored and taken into account by controllers (intel_corporation_intel_2023, 46). Monitoring features therefore play a crucial role in effective feature utilization as they allow controllers to observe changes in metrics resulting from a tuning decision.
@@ 3. Review Motivation
By observation of published literature, it is clear that QoS enforcement features are conspicuous by their absence in general-purpose cloud computing resource management surveys, see Table 1.
@@
(mittal_survey_2017, 65)
(gracioli_survey_2015, 35, 63)
(kawashima_vision_2021, 52)
(hameed_survey_2016, 37, 59, 43, 9, 39, 13, 11, 44)
@@
Cloud environments rarely see hard execution time deadlines on tasks, as is common in real-time systems, but rather statements of performance intent in the form of service level objectives (SLOs). Resource management optimization for cloud environments is therefore typically discussed in a broader context and at a higher level of abstraction than real-time systems. Thus, lower-level sources of congestion such as LLC and memory bandwidth are often overlooked, excluded, or disregarded in cloud resource management surveys (hameed_survey_2016, 37, 59, 43, 9, 39, 13, 11, 44). Even so, increased resource efficiency and task consolidation achieved through QoS enforcement feature utilization is not exclusive to real-time systems.
We thus recognize the need for a broad review of literature approaching shared resource management using hardware-level QoS enforcement features to gain insight into past and current trends in research. This review serves as a basis for understanding how the academic community approaches resource management issues using hardware-implemented QoS enforcement features. By understanding the current approaches and their traits, we aim to detail the major challenges that hinder the practicality and widespread adoption of QoS enforcement in the cloud.
@@ 3.1. Related Surveys
A few systematic reviews have previously been conducted that cover neighboring and overlapping areas of research. In their 2015 article, (gracioli_survey_2015, 35) (gracioli_survey_2015, 35) surveyed cache management approaches for embedded real-time systems. They recognize that server processors supporting Intel CAT, the only commercially available feature at the time, are not targeted at the embedded systems market surveyed in their paper. However, they still make the case for cache allocation’s usefulness for the optimization of soft real-time applications.
@@
Publications with a focus on resource management in the cloud often approach the problem at a higher abstraction layer than papers with lower-level performance as their context, such as those from the real-time computing community. This often means that CPU cache and memory bandwidth management on the node level is out of scope. Such is the case in the surveys by (hameed_survey_2016, 37) (hameed_survey_2016, 37), (liaqat_federated_2017, 59) (liaqat_federated_2017, 59), and (hong_resource_2019, 43) (hong_resource_2019, 43).
@@
With a clear focus on cache partitioning, (mittal_survey_2017, 65) (mittal_survey_2017, 65) surveyed techniques for mitigating the increased shared cache congestion that follows from increasing core counts in processors. This includes early publications utilizing CAT, as well as a large number of software-based approaches using techniques such as page coloring.
@@
(bermejo_virtual_2020, 9) (bermejo_virtual_2020, 9), (helali_survey_2021, 39) (helali_survey_2021, 39), and (chaurasia_comprehensive_2021, 13) (chaurasia_comprehensive_2021, 13) surveyed techniques of workload consolidation in the cloud. The first with a specific focus on virtual machine (VM) consolidation. (kawashima_vision_2021, 52) (kawashima_vision_2021, 52) acknowledge the usefulness of cache partitioning when optimizing the performance of virtual network functions (VNFs) that power much of the networking in modern clouds.
@@
(lugo_survey_2022, 63) (lugo_survey_2022, 63) surveyed works on process interference for real-time systems published between 2015 and 2020. The works included proposed techniques that reduce congestion in LLC, main memory, as well as in the main memory bus. A few of the works utilize QoS enforcement features, while others use software-based approaches.
@@
With the rise of Kubernetes as the de-facto standard container orchestrator, (carrion_kubernetes_2022, 11) (carrion_kubernetes_2022, 11) identifies the need to survey the current approaches and ongoing issues in Kubernetes scheduling. They find that Kubernetes is used in many different environments, running different types of workloads on different types of hardware, which dictates the type of scheduling approaches leveraged.
@@
In a recent publication, (huang_survey_2023, 44) (huang_survey_2023, 44) review works on resource management in the context of cloud-native mobile computing. This has received an increasing amount of attention recently due to the architecture of fifth and sixth-generation mobile communication networks (5G and 6G) being based on cloud-native technology. Most of the works reviewed focus on the management of networking resources through the utilization of network slicing or network virtualization.
@@ 4.1. Search Strategy
Our search strategy is based on the reference snowballing process outlined by (wohlin_guidelines_2014, 106) (wohlin_guidelines_2014, 106) and follows a simple algorithm. First, a starting set of core publications is identified by querying a digital literature database using a defined set of search terms and inclusion criteria. Then, both backward and forward snowballing is used to identify additional papers that fit the inclusion criteria. The snowballing process is repeated after identified papers are added to the set of included papers. This process is repeated until no additional papers that fit the inclusion criteria are found. This process is outlined in Figure 5.
Figure 5. Reference snowballing search strategy used.:
@@
Our motivations for selecting this approach instead of a more traditional database search approach (i.e. (kitchenham_evidence-based_2004, 56) (kitchenham_evidence-based_2004, 56)) are as follows: We survey how the academic community leverages a defined set of hardware features in research. We can thus find our starting set of articles by explicitly searching for the names of the features. This should prove a good basis for snowballing as authors typically cite a small set of early papers as sources for further details of a feature. Such publications become natural “roots” in our reference tree. Moreover, it was shown by (badampudi_experiences_2015, 6) (badampudi_experiences_2015, 6) that snowballing is comparable, if not more reliable, than a traditional database search approach.
@@ 4.2. Threats to Review Validity
The snowballing process is highly dependent on the quality of the starting set of papers used to seed the search. A starting set that is too small, or not diverse enough can leave “dark spots” where many papers are left out. One reason for this is a same author bias that may emerge. As researchers often cite their previous work due to relevance, a snowballing search may become heavily biased towards a few researchers’ works (jalali_systematic_2012, 50). The same phenomenon may be true for publishers instead of individual researchers (wohlin_guidelines_2014, 106). However, the nature of our review and its focus on a select set of hardware features should limit the impact of this bias, as researchers reference a set of early papers when discussing these features.
@@ 4.3. Inclusion Criteria
No starting date of the publication time span is defined as hardware QoS features have only recently become available in commodity off-the-shelf server processors. Intel’s CAT was originally introduced in the Xeon E5-2600 v3 family of processors (herdrich_cache_2016, 41) which launched in the third quarter of 2014. Thus, 2014 serves as a starting date for CAT publications.
We focus our review on hardware-based control mechanisms, despite the long-standing existence of software-based QoS enforcement approaches. Software-based methods, such as page coloring (taylor_tlb_1990, 99) or MemGuard (yun_memguard_2013, 118), offer similar functionalities to their hardware counterparts and have likely influenced the design of current hardware features.
While processor-agnostic software control strategies have their own advantages and use cases, potentially addressing some limitations of current hardware mechanisms, chip manufacturers continue to enhance hardware-level control mechanisms with each new generation (sohal_closer_2022, 95). This ongoing improvement suggests that the future of QoS enforcement lies in hardware-level mechanisms, which require lower-level control and platform-specific optimizations for optimal performance. Therefore, we concentrate on hardware-level control mechanisms in this review.
@@ 4.4. Starting Set Identification
Google Scholar [^3] http://scholar.google.com was used as a search engine when querying for potential starting set articles. This had the advantage of minimizing potential publisher bias as the search covers a large number of publishing sources. A separate search was conducted for each of the hardware features of interest. The exact search terms used were:
@@
A set of seed publications were selected from these lists of publications based on our inclusion criteria, publisher diversity, and relevance. This start set of papers are listed in Table 2 and a graph of their reference relationships may be seen in Figure 6. While this starting set is small and may contain important gaps in the literature, it is a large and diverse enough starting point. The snowballing process will help fill in these gaps as the included papers reference, or are referenced by, other important publications.
@@
(funaro_ginseng_2016, 31) 1 (funaro_ginseng_2016, 31)
(xu_vcat_2017, 113) 1 (xu_vcat_2017, 113)
(selfa_application_2017, 92) 1 (selfa_application_2017, 92)
(xu_dcat_2018, 111) 1 (xu_dcat_2018, 111)
(pons_improving_2018, 80) 1 (pons_improving_2018, 80)
(park_hypart_2018, 75) 2 (park_hypart_2018, 75)
(xiang_dcaps_2018, 108) 1 (xiang_dcaps_2018, 108)
(kim_application_2019, 55) 1 (kim_application_2019, 55)
(park_copart_2019, 74) 2 (park_copart_2019, 74)
(xiang_emba_2019, 109) 2 (xiang_emba_2019, 109)
(zhang_libra_2021, 122) 3 (zhang_libra_2021, 122)
(sohal_closer_2022, 95) 1 (sohal_closer_2022, 95)
(farina_assessing_2022, 27) 2 (farina_assessing_2022, 27)
@@
Figure 6. Arc diagram illustrating the reference relationships between the papers selected for the starter set. Publications are chronologically ordered from top to bottom. The size of the nodes corresponds to the number of connections they have.:
@@ 4.5. Reference Snowballing
Using the snowballing search method, we iterated upon the defined starting set. At the time of the final snowballing algorithm iteration, the search set had increased to $3685$ references. Out of these $3685$ publications, $78$ papers were identified to meet the inclusion criteria of our review. Figure 7 illustrates the included articles’ publication dates and how interest from researchers in both academia and industry has evolved over time.
@@
Figure 7. Distribution of articles utilizing platform features grouped by their year of publication. Papers are classified as academic when all authors are affiliated with academic institutions, whereas cooperative publications have one or more authors affiliated with a private company. Industry papers have no affiliation with an academic institution. (*2023 is limited to the first six months of the year.):
@@ 5. Taxonomy
Figure 8. Taxonomy organization.:
@@ 5.1. Feature Evaluations
The function of QoS enforcement features implemented in hardware has been experimentally analyzed by a number of different works (sohal_closer_2022, 95, 122, 41, 27, 45, 91, 54, 103, 30). It is important to verify the behavior of hardware features so that they may be used as a basis for other research. However, the functionality may change between chip generations due to architectural differences (sohal_closer_2022, 95). This further highlights the importance of analyzing the behavior of new implementations as the features mature over time. Additionally, some works (bechtel_denial–service_2022, 8, 86, 29) evaluate if QoS enforcement features may be used as solutions to problems in other research areas. Other publications (shang_coplace_2021, 94) use QoS enforcement features to demonstrate problems and subsequent solutions to resource management problems.
@@
(herdrich_cache_2016, 41) Introduction and feature showcase of CAT.
(sohal_closer_2022, 95) Experimental evaluation of CAT and MBA.
(zhang_libra_2021, 122) Highlights the shortcomings of MBA.
(farina_assessing_2022, 27) Experimental evaluation of MBA for RT systems.
(ihre_sherif_evaluation_2021, 45) Experimental evaluation of CAT.
(schramm_impact_2019, 91) Experimental evaluation of CAT for VNF systems.
(qin_when_2018, 84) Study of cache partitioning in correlation to workload cache characteristics.
(kim_simple_2017, 54) Evaluation of cache partitioning as a tool to meet workload latency targets.
(veitch_performance_2017, 103) Experimental evaluation of CAT for VNF systems.
(castiglione_optimal_2018, 30) Highlights the potential usefulness of CAT for security purposes.
@@
(qiu_is_2021, 86) Evaluation of FaaS for LC workload hosting.
(bechtel_denial–service_2022, 8) Evaluation of LLC partitioning as DoS protection on iGPU platforms.
(farshin_reexamining_2020, 29) Evaluation of Intel Data Direct I/O.
(shang_coplace_2021, 94) Intelligent Guest to Host Page Mapping.